OpenStudy (anonymous):

I am really hoping to get help with solving a radical expression nroot{3}(125(x+8)^3

OpenStudy (anonymous):

I'm slightly confused as to what the problem is.

OpenStudy (anonymous):

\[\sqrt[3]{125(x+8)^{3}}\] I am trying to solve this radical expression

OpenStudy (helder_edwin):

\[ \large \sqrt[3]{125(x+8)^3}=\sqrt[3]{5^3(x+8)^3}=5(x+8) \]